Based on a​ poll, among adults who regret getting​ tattoos, 12​% say that they were too young when they got their tattoos. Assume that eight adults who regret getting tattoos are randomly​ selected, and find the indicated probability. Complete parts​ (a) through​ (d) below.


1
Expert's answer
2021-10-07T11:55:25-0400

Based on a poll, among adults who regret getting tattoos, 12% say that they were too young when they got their tattoos. Assume that eight adults who regret getting tattoos are randomly selected, and find the indicated probability. Complete parts (a) through (d) below.

a. Find the probability that none of the selected adults say that they were too young to get tattoos. (Round to four decimal places as needed.)

b. Find the probability that exactly one of the selected adults says that he or she was too young to get tattoos. (Round to four decimal places as needed.)

c. Find the probability that the number of selected adults saying they were too young is 0 or 1. (Round to four decimal places as needed.)

d. If we randomly select eight adults, is 1 a significantly low number who say that they were too young to get tattoos? because the probability that of the selected adults say that they were too young is 0.05.

"P(X=x)= C^n_xp^x(1-p)^{n-x}"

a. "P(X=0) = C^8_0 \times 0.12^0 \times 0.88^8 = 0.3596"

b. "P(X=1) = C^8_1 \times 0.12^1 \times 0.88^7 = 0.3923"

c. "P(X=0) + P(X=1) = 0.3596 + 0.3923 = 0.7519"

d. No, as probability that at most one of the selected adult lay that they were too young is greater then 0.05.